# Dynamical phase transitions in the non-reciprocal Ising model

### v1

@@ X Coarsening dynamics: scroll waves and planar waves
We now qualitatively discuss the way in which the system coarsens into an ordered state in 3D. In the simulations used to calculate the order parameters $R$ and $\mathcal{L}$ (Figs. 12, 13,14, and 20), the spins were initialized in an ordered initial condition (all spins up), making the transition to steady state in the swap and static-order phases relatively fast. When fully disordered initial conditions are used (each spin is set to $+1$ or $-1$ randomly), however, the system can coarsen into long-lived metastable states, before reaching steady state. To show this, we present in Fig. 29a the same plot as Fig. 13a ($R$ vs. $\tilde{J}$ in 3D) but with random initial conditions, and average taken over a single simulation running for a very long time. In both the swap phase ($1.4\lesssim\tilde{J}\lesssim 2.3$) and static order phase ($\tilde{J}\gtrsim 2.3$) we see that for some of the realizations the system reaches the same steady state as with ordered initialization (shown in semi-transparent lines). However, sometimes the measured $R$ is much lower than the one obtained with ordered initialization, suggesting that the system got trapped in a metastable state. We cannot exclude that several equally stable states exist.
@@
The coarsening dynamics is most easily understood by following the defect filaments of the scroll waves. The defects are found using the size dependent synchronization order parameter, $s(l)$, in Eq. (62). Note that $s(l)$ is 1 if all the sites in the box are identical, and it is close to zero if the box has a mixture of sites in all four possible states. Thus, we classify defects as sites on the coarse-grained grid in which $s(l)$ is smaller than a threshold $s_{\rm max}$. The box size $l$ and the $s$-threshold are fine-tuned to detect the defect filaments while filtering out noise (in Fig. 29 $l=3$ and $s_{\rm max}=0.2$).
Figure. 29e shows an example of the dynamics of the defect filaments. At the beginning the system coarsens into a complex mixture of defect lines. After a while, only two defect lines survive. The lines are connected onto themselves by virtue of the periodic boundary conditions, having the topology of rings on a 3-dimensional torus with opposite winding numbers ($1$ and $-1$). Thus, they cannot shrink by themselves to a point. Finally, the two filaments collide, forming a new ring with zero winding number that shrinks and disappears. The system then stays in a planar waves state for a very long time. Movie 3 shows another example of the filament dynamics that ends in a defect-free swap phase.
The occurrences in which the system escapes scroll-waves state into swap or static order, without any observation of transitions from the swap or static order phase into scroll waves, suggest that these are metastable states. This is not the case, however, in the high $\tilde{K}$ limit, where the scroll waves appear to be stable, as explained in Sec. VIII.3.
@@ XI Schematic phase diagram in the fully anti-symmetric case
In 2D, our numerical results suggest that both the swap phase and the static order phase are unstable, as long as $K\neq 0$. This leads to the conclusion that there is a single stable phase in the thermodynamic limit: disorder (see Fig. 30a). Note that the same conclusion applies to the 1D case, as is shown in Appendix D.
In 3D (see Fig. 30b), our numerical results suggest that static order is unstable but that the swap phase is stable, with a second-order-like phase transition from disorder to swap featuring 3D XY critical exponents. Regions with high $\tilde{J}$ or $\tilde{K}$ have not been revealed in full: in the large $\tilde{J}$ limit, synchronization decreases, possibly to zero, due to the difficulty of droplet-induced swap to synchronize. In the high $\tilde{K}$ regime, scroll waves completely destabilize the long-range order. Both mechanisms are capable of destroying a large portion of the stable swap phase regime, and there is also a possibility that they would meet in the high $\tilde{J}$ and high $\tilde{K}$ regime, leading to a compact swap phase in $(\tilde{J},\tilde{K})$ space rather than an infinite one. Due to the limitations of our simulations, these interesting possibilities lie beyond the scope of this work, which has focused mostly on the behavior at and close to the Hopf bifurcation, where the swap phase appears to be stable and critical exponents can be calculated.
Fig. 30: Schematic phase diagram of the fully anti-symmetric case. (a) Schematic 2D phase diagram and (b) schematic 3D phase diagram. To be compared with the MF prediction, Fig. 4. Green line is second-order-like phase transition with critical exponents compatible with those of the 3D XY model. Grey dashed lines are phase transitions that have not been revealed in full and are possibly first-order-like.
@@ XII General asymmetric couplings
In the previous sections, we made two assumptions. First, we assumed that the interaction between the two spin species is fully anti-symmetric, i.e. $K_{+}=0$ in Eq. 12. Second, we assumed that both species have the same nearest-neighbors interaction strength $J$. These two conjoint assumptions lead to a symmetry between the two species which renders the model $C_{4}$ symmetric (instead of the less restrictive $\mathbb{Z}_{2}$ symmetry with only up-down symmetry). Therefore, we need to test whether our results rely on this symmetry or are robust to its breakdown.
In this section, we break the $C_{4}$ symmetry by considering a scenario where $K_{+}\neq 0$.
@@ XII.1 Mean-field analysis
Expanding Eq. (23) in the small $m_{\alpha}$ and $\nabla m_{\alpha}$ limit, we obtain,
@@
$-\frac{1}{3}\left(\tilde{J}m_{\alpha}+\tilde{K}_{\alpha\beta}m_{\beta}\right)^{3}.$
Taking $D=0$ (as was done in Sec. V.2 for the fully anti-symmetric case), the eigenvalues $\lambda_{1}$ and $\lambda_{2}$ of the Jacobian matrix $\mathcal{J}$ at $m_{A}=m_{B}=0$ are
@@
Thus, upon changing $\tilde{J}$, the possible phases and bifurcations depend on the value of $\tilde{K}_{+}^{2}-\tilde{K}_{-}^{2}$, as shown in Fig. 31 for the the full homogeneous MF equation, Eq. (23) with $D=0$.
@@
For $\tilde{K}_{+}^{2}-\tilde{K}_{-}^{2}<0$ ($\tilde{K}_{+}<\tilde{K}_{-}$) the system is disordered when $\tilde{J}<1$, and has a Hopf bifurcation at $\tilde{J}=1$, taking the system from disorder to a limit cycle (Fig. 31a). Note that unlike the $K_{+}=0$ case studied throughout most of the paper, here the limit cycle is not rotationally-symmetric even in the $\tilde{J}-1\ll 0$ limit. This is easily seen when writing the homogeneous MF equation in polar coordinate
@@
$\begin{split}\partial_{t}r&=-\left(1-\tilde{J}-\tilde{K}_{+}\sin(2\theta)\right)r+\mathcal{O}\left(r^{3}\right)\\ \partial_{t}\theta&=-\tilde{K}_{-}+\tilde{K}_{+}\cos(2\theta)+\mathcal{O}\left(r^{2}\right)\end{split}$ (65)
@@
where the leading order term in $r$ is $\theta$-dependent and a separation between the time scales of $r$ and $\theta$ can not be invoked as in Sec. V.2. Instead, the limit cycle is elliptical close to the Hopf bifurcation [151].
@@
Fig. 31: Bifurcations in the general asymmetric case. Flow in $\left(m_{A},m_{B}\right)$ space calculated from the full MF equation, Eq. (23) with $D=0$ and $\tilde{K}_{-}=0.2$, and with (a) $\tilde{K}_{+}=0.1$, (b) $\tilde{K}_{+}=0.3$, and (c) $\tilde{K}_{+}=1.1$. In each panel, $\tilde{J}$ is increased from left to right to capture the different phases. In (a) $\tilde{J}=0.4,1.2,1.5,1.7$, in (b) $\tilde{J}=0.4,0.9,1.4,2$ and in (c) $\tilde{J}=0.1,2.3,3.1$. The arrows mark the bifurcations that connect every two configurations: Hopf, saddle-node on an invariant circle (SNIC), saddle-node (SN) or pitchfork (PF). The bifurcations are consistent with the analysis of the expanded MF equation (Sec. XII.1).
@@
Increasing $\tilde{J}$ leads to additional bifurcations (Fig. 31a): first, a SNIC bifurcation where the limit cycle is destroyed by the emergence of two (rather than 4 for the $K_{+}=0$ case) pairs of stable and saddle points, and for an even larger $\tilde{J}$, a saddle-node bifurcation where two additional pairs of stable and saddle points emerge. We note that this behavior is observed for $\tilde{K}_{+}$ and $\tilde{K}_{-}$ much smaller than $1$ ($0.1$ and $0.2$ respectively in the Fig. 31). For larger values, a saddle-node bifurcation can occur such that the ferromagnetic phase coexists with the limit cycle (this scenario is explored in  [143]).
@@
For $10\leq\tilde{K}_{+}^{2}-\tilde{K}_{-}^{2}<$ 1 the system is disordered when $\tilde{J}<1-\sqrt{\tilde{K}_{+}^{2}-\tilde{K}_{-}^{2}}$ and has a pitchfork bifurcation taking the system from disorder to static order with two stable fixed points. In that case, there is no oscillatory phase (Fig. 31b). Finally, for $\tilde{K}_{+}^{2}-\tilde{K}_{-}^{2}>1$ there are two stable fixed points already at $\tilde{J}=0$ (Fig. 31c) and thus the only stable homogeneous phase is the static order phase (although $\tilde{J}$ must be larger than zero for coarse-graining to take place). We note that the absence of oscillations for $\tilde{K}_{+}>\tilde{K}_{-}$ is due to the fact that the two species, while having non-reciprocal interactions, “agree” with each other on whether they like to align or anti-align, making a magnetization flip unfavorable for either species.
@@
Figure 32 shows a cut through the three-dimensional (depending on $\tilde{J},\tilde{K}_{+}$, and $,\tilde{K}_{-}$) phase diagram, for a fixed $\tilde{K}_{+}$ (set to be $0.3$). The main difference between this phase diagram and the one in Fig. 4 showing the $\tilde{K}_{+}=0$ case, is the pitchfork bifurcation line from a paramagnetic to ferromagnetic phase in the $\tilde{K}_{-}\leq\tilde{K}_{+}$ region, which is merely a point in Fig. 4.
@@
Fig. 32: Mean-field phase diagram of the general asymmetric case. The diagram is shown as a function of $\tilde{J}$ and $\tilde{K}_{-}$. Reciprocal inter-species interaction coupling is set to be $\tilde{K}_{+}=0.3$. The lines color code is the same as Fig. 4. Obtained from numerical analysis of Eq. (23) with $D=0$.
@@ XII.2 Swap phase
We perform MC simulations in both 2D and 3D for the general asymmetric case, i.e. with the inclusion of reciprocal inter-species interactions. In 2D, we find that the swap phase is unstable, as in the case of fully anti-symmetric interactions (Sec. VII). This is shown in Fig. 33 in which $R$ and $\mathcal{L}$ are plotted as a function of $\tilde{J}$ with fixed $\tilde{K}_{+}$ and $\tilde{K}_{-}$. In the region in which $\mathcal{L}\neq 0$ (indicating the swap phase), both $R$ and $\mathcal{L}$ decrease with system size. In appendix E we show color maps of $R$ and $\mathcal{L}$ as a function of $\tilde{J}$ and $\tilde{K}_{-}$ with fixed non-zero $\tilde{K}_{+}$. The color maps show that, similarly to the $\tilde{K}_{+}=0$ case in Fig. 9, $R$ and $\mathcal{L}$ diminish as system size increases.
@@
In 3D, MC simulations suggest that a stable swap phase still exists, even when reciprocal interactions are added. Figure 34a-b shows color maps of $R$ and $\mathcal{L}$ for a fixed $\tilde{K}_{+}$ and $L=80$. Figure 34c-d then analyzes a cut through the color map for fixed $\tilde{K}_{-}$. As $\tilde{J}$ is varied, a phase transition from disorder to swap is observed and becomes more sharp as $L$ increases. The swap phase is stable for $\tilde{J}\lesssim 1.7$ but destabilizes at higher values, as is discussed next.
@@
Fig. 34: Stable swap phase in 3D in the general asymmetric case. (a) $R$ and (b) $\mathcal{L}$ color map as a function of $\tilde{J}$ and $\tilde{K}_{-}$ with $\tilde{K}_{+}=0.1$ and $L=80$. Dashed grey line shows the fixed $\tilde{K}_{-}$ cut analyzed in panels b and c. (c) $R$ and (d) $\mathcal{L}$ as a function $\tilde{J}$ with $\tilde{K}_{+}=0.1$ and $\tilde{K}_{-}=0.3$, for different system sizes. Inset in c shows a snapshot of $\theta$ in a 2D-cut of the 3D system, with $\tilde{J}=2$, $L=80$, and the same $\tilde{K}_{+}$ and $\tilde{K}_{-}$ values as in the main panel. Obtained from MC simulations.

### v2

@@ X Coarsening dynamics: scroll waves and planar waves
We now qualitatively discuss the way in which the system coarsens into an ordered state in 3D. In the simulations used to calculate the order parameters $R$ and $\mathcal{L}$ (Figs. 11, 12 and 18), the spins were initialized in an ordered initial condition (all spins up), making the transition to steady state in the swap and static-order phases relatively fast. When fully disordered initial conditions are used (each spin is set to $+1$ or $-1$ randomly), however, the system can coarsen into long-lived metastable states, before reaching steady state. To show this, we present in Fig. 27a the same plot as Fig. 12a ($R$ vs. $\tilde{J}$ in 3D) but with random initial conditions, and average taken over a single simulation running for a very long time. In both the swap phase ($1.4\lesssim\tilde{J}\lesssim 2.3$) and static-order phase ($\tilde{J}\gtrsim 2.3$) we see that for some of the realizations the system reaches the same steady state as with ordered initialization (shown in semi-transparent lines). However, sometimes the measured $R$ is much lower than the one obtained with ordered initialization, suggesting that the system got trapped in a metastable state.
@@
The coarsening dynamics is most easily understood by following the defect filaments of the scroll waves (Fig. 27e). The defects are found using the size-dependent synchronization parameter, $s(l)$, defined in Eq. (71). Note that $s(l)$ is 1 if all the sites in the box are identical, and it is close to zero if the box has a mixture of sites in all four possible states. Thus, we classify defects as sites on the coarse-grained grid in which $s(l)$ is smaller than a threshold $s_{\rm max}$. The box size $l$ and the $s$-threshold are fine-tuned to detect the defect filaments while filtering out noise ($l=3$ and $s_{\rm max}=0.2$ in Fig. 27e).
In the example shown in Fig. 27e, the system coarsens into a complex mixture of defect lines shortly after initialization with disordered initial conditions. After a while, only two defect lines survive. The lines are connected onto themselves by virtue of the periodic boundary conditions, having the topology of rings on a 3-dimensional torus with opposite winding numbers ($1$ and $-1$). Thus, they cannot shrink by themselves to a point. Finally, the two filaments collide, forming a new ring with zero winding number that shrinks and disappears. The system then stays in a planar waves state for a very long time. Movie 3 shows another example of the filament dynamics that ends in a defect-free swap phase.
The occurrences in which the system escapes scroll-waves state into swap or static order, without any observation of transitions from the swap or static-order phase into scroll waves, suggest that these are metastable states. This is not the case, however, in the high $\tilde{K}$ limit, where the scroll waves appear to be stable, as explained in Sec. VIII.3.
@@ XI Schematic phase diagram in the fully anti-symmetric case
The results from Secs. VIIVIII, and IX enable us to construct a schematic phase diagram in both 2D and 3D for the fully anti-symmetric case.
In 2D, our numerical results suggest that both the swap phase and the static-order phase are unstable, as long as $K\neq 0$. This leads to the conclusion that there is a single stable phase in the thermodynamic limit: disorder (see Fig. 28a). The same conclusion applies to the 1D case, as is shown in Appendix B.
In 3D (see Fig. 28b), our numerical results suggest that static order is unstable but that the swap phase is stable, with a second-order-like phase transition from disorder to swap featuring 3D XY critical exponents. Regions with high $\tilde{J}$ or $\tilde{K}$ have not been revealed in full: in the large $\tilde{J}$ limit, synchronization decreases, possibly to zero, due to the difficulty of droplet-induced swap to synchronize. In the high $\tilde{K}$ regime, scroll waves completely destabilize the long-range order. Both mechanisms are capable of destroying a large portion of the stable swap phase regime, and there is also a possibility that they would meet in the high $\tilde{J}$ and high $\tilde{K}$ regime, leading to a compact swap phase in $(\tilde{J},\tilde{K})$ space rather than an infinite one. Due to the limitations of our simulations, these interesting possibilities lie beyond the scope of this work, which has focused mostly on the low $\tilde{J}$ and $\tilde{K}$ regime, where the swap phase appears to be stable and critical exponents can be calculated.
Fig. 28: Schematic phase diagram of the fully anti-symmetric case. (a) Schematic 2D phase diagram and (b) schematic 3D phase diagram. To be compared with the mean-field prediction, Fig. 4. The green line is the second-order-like phase transition with critical exponents compatible with those of the 3D XY model. Grey dashed lines are phase transitions that have not been revealed in full and are possibly first-order-like. In the striped area, it is not clear from numerical simulations whether the system is in the swap or disordered phase.
@@ XII General asymmetric couplings
In the previous sections, we have focused on the fully anti-symmetric case where $K_{+}=0$ in Eq. (12), for which our model has a $C_{4}$ symmetry. In this section, we test whether our results rely on this additional symmetry by considering $K_{+}\neq 0$, for which the symmetry is reduced to the diagonal $\mathbb{Z}_{2}$ described in Eq. (24).
@@ XII.1 Mean-field analysis
Expanding Eq. (23) as was done for the fully anti-symmetric case (Sec. V.2), we obtain,
@@
$-\frac{1}{3}\left(\tilde{J}m_{\alpha}+\tilde{K}_{\alpha\beta}m_{\beta}\right)^{3}$
where $\tilde{K}_{\alpha\beta}\equiv K_{\alpha\beta}/(k_{B}T)$ and $K_{\alpha\beta}$ is given by Eq. (12). Assuming a spatially homogeneous state, the eigenvalues $\lambda_{1}$ and $\lambda_{2}$ of the Jacobian matrix $\mathcal{J}$ at $m_{A}=m_{B}=0$ are
@@
Thus, upon changing $\tilde{J}$, the possible phases and bifurcations depend on the value of $\tilde{K}_{+}^{2}-\tilde{K}_{-}^{2}$. This is demonstrated in Fig. 29, which shows phase portraits of the spatially homogeneous full mean-field equation, Eq. (23), for different $\tilde{K}_{+}^{2}-\tilde{K}_{-}^{2}$ values.
@@
For $\tilde{K}_{+}^{2}-\tilde{K}_{-}^{2}<0$ (which is equivalent to $\tilde{K}_{+}<\tilde{K}_{-}$ when both $\tilde{K}_{+}$ and $\tilde{K}_{-}$ are assumed to be positive), Fig. 29a, the system is disordered when $\tilde{J}<1$, and undergoes a Hopf bifurcation at $\tilde{J}=1$, transitioning from disorder to a limit cycle. Unlike the $K_{+}=0$ case studied throughout most of the paper, here the limit cycle is not rotationally-symmetric even in the $\tilde{J}-1\ll 1$ limit. This is seen when writing the spatially homogeneous mean-field equation in polar coordinates,
@@
$\begin{split}\partial_{t}r&=-\left(1-\tilde{J}-\tilde{K}_{+}\sin(2\theta)\right)r+\mathcal{O}\left(r^{3}\right)\\ \partial_{t}\theta&=-\tilde{K}_{-}+\tilde{K}_{+}\cos(2\theta)+\mathcal{O}\left(r^{2}\right).\end{split}$ (74)
@@
The leading order term in $r$ is $\theta$-dependent, and does not simply scale with $\tilde{J}-1$. Hence, a separation between the time scales of $r$ and $\theta$ cannot be invoked as in Sec. V.2. Instead, the limit cycle is elliptical near the Hopf bifurcation [85].
@@
Fig. 29: Bifurcations in the general asymmetric case. Flow in $\left(m_{A},m_{B}\right)$ space calculated from the spatially homogeneous full mean-field equation, Eq. (23), for $\tilde{K}_{-}=0.2$, and (a) $\tilde{K}_{+}=0.1$, (b) $\tilde{K}_{+}=0.3$, and (c) $\tilde{K}_{+}=1.1$. In each panel, $\tilde{J}$ is increased from left to right to capture the different phase portraits. In (a) $\tilde{J}=0.4,1.2,1.5,1.7$, in (b) $\tilde{J}=0.4,0.9,1.4,2$ and in (c) $\tilde{J}=0.1,2.3,3.1$. The arrows mark the bifurcations that connect every two configurations: Hopf, saddle-node on an invariant circle (SNIC), saddle-node (SN) or pitchfork (PF).
@@
Increasing $\tilde{J}$ leads to additional bifurcations. First, a ($\mathbb{Z}_{2}$-symmetric) SNIC bifurcation where the limit cycle is destroyed by the emergence of two (rather than four for the $K_{+}=0$ case) pairs of stable and saddle points. The stable points that emerge lie in the first and third quadrants of the $(m_{A},m_{B})$ plane, corresponding to states where the two species are aligned. Then, for even larger values of $\tilde{J}$, a saddle-node bifurcation occurs where two additional pairs of stable and saddle points emerge, yet static order is maintained. The new stable points lie in the second and fourth quadrants, representing anti-aligned states. This behavior is observed for $\tilde{K}_{+}$ and $\tilde{K}_{-}$ much smaller than $1$ ($0.1$ and $0.2$ respectively in the Fig. 29). For larger values, a saddle-node bifurcation can occur such that the ferromagnetic phase coexists with the limit cycle (this scenario is explored in Ref. [151]).
@@
For $10\leq\tilde{K}_{+}^{2}-\tilde{K}_{-}^{2}<$ 1 (Fig. 29b), the system is disordered when $\tilde{J}<1-\sqrt{\tilde{K}_{+}^{2}-\tilde{K}_{-}^{2}}$ and undergoes a pitchfork bifurcation at $\tilde{J}=1-\sqrt{\tilde{K}_{+}^{2}-\tilde{K}_{-}^{2}}$ transitioning from disorder to static order with two stable fixed points of aligned states. In that case, there is no oscillatory phase. When $\tilde{J}=1+\sqrt{\tilde{K}_{+}^{2}-\tilde{K}_{-}^{2}}$, an additional pitchfork bifurcation occurs where two saddle points emerge from the origin $(0,0)$, which transitions from a saddle to an unstable point. As $\tilde{J}$ increases further, a saddle-node bifurcation occurs where two pairs of anti-aligned stable and saddle points appear. These two additional bifurcations maintain the system in a static-order phase. In the $\tilde{K}_{+}^{2}-\tilde{K}_{-}^{2}>1$ case (Fig. 29c), the bifurcations are the same as in the $10\leq\tilde{K}_{+}^{2}-\tilde{K}_{-}^{2}<$ 1 case, except that there are already two stable fixed points at $\tilde{J}=0$. As a result, the system remains in the static-order phase for any value of $\tilde{J}$ (although $\tilde{J}$ must greater than zero for coarse-graining to take place). The absence of oscillations for $\tilde{K}_{+}>\tilde{K}_{-}$ arises because, despite having nonreciprocal interactions, the two species “agree” on whether they prefer to align or anti-align. This makes a magnetization flip unfavorable for either species.
@@
Figure 30 shows a slice through the three-dimensional phase diagram (which depends on $\tilde{J},\tilde{K}_{+}$, and $,\tilde{K}_{-}$), with $\tilde{K}_{+}$ fixed at $0.3$. The main difference between this phase diagram and the one in Fig. 4 (which shows the $\tilde{K}_{+}=0$ case) is the pitchfork bifurcation line separating the paramagnetic and ferromagnetic phases in the $\tilde{K}_{-}\leq\tilde{K}_{+}$ region, which is just a point in Fig. 4.
@@
Fig. 30: Mean-field phase diagram of the general asymmetric case. The diagram is shown as a function of $\tilde{J}$ and $\tilde{K}_{-}$, with the reciprocal inter-species interaction coupling set to $\tilde{K}_{+}=0.3$. The saddle-node bifurcations illustrated in Fig. 29 are not drawn in this plot. Obtained from numerical analysis of the spatially homogeneous Eq. (23).
@@ XII.2 Swap phase
We perform Monte-Carlo simulations in both 2D and 3D for the general asymmetric case, i.e. with the inclusion of reciprocal inter-species interactions. In 2D, we find that the swap phase is unstable, as in the case of fully anti-symmetric interactions (Sec. VII). This is shown in Fig. 31 in which $R$ and $\mathcal{L}$ are plotted as a function of $\tilde{J}$ with fixed $\tilde{K}_{+}$ and $\tilde{K}_{-}$ (a-b) and as a function of $\tilde{K}_{+}$ with fixed $\tilde{J}$ and $\tilde{K}_{-}$ (c-d). In the region in which $\mathcal{L}\neq 0$ (indicating the swap phase), both $R$ and $\mathcal{L}$ decrease with system size. Furthermore, in appendix F we show color maps of $R$ and $\mathcal{L}$ as a function of $\tilde{J}$ and $\tilde{K}_{-}$ with a fixed non-zero value of $\tilde{K}_{+}$. These color maps show that, similarly to the $\tilde{K}_{+}=0$ case shown in Fig. 9a-d, $R$ and $\mathcal{L}$ diminish as system size increases in the swap regime.
@@
In 3D, Monte-Carlo simulations suggest that a stable swap phase still exists, even when reciprocal interactions are added. Figure 32 shows color maps of $R$ and $\mathcal{L}$ for a fixed $\tilde{K}_{+}$ and $L=80$. A region with non-zero $R$ and $\mathcal{L}$, corresponding to a swap phase, is observed. Figure 33 then analyzes a line cut through the color maps for a fixed value of $\tilde{K}_{-}$. As $\tilde{J}$ is varied, a continuous phase transition from disorder to swap is observed (at $\tilde{J}\approx 1.45$) and becomes more sharp as $L$ increases. The swap phase is stable for $\tilde{J}\lesssim 1.7$ but destabilizes at higher values, as is discussed next.
@@
Fig. 32: Order parameters in 3D in the general asymmetric case. (a) $R$ and (b) $\mathcal{L}$ color map as a function of $\tilde{J}$ and $\tilde{K}_{-}$ with $\tilde{K}_{+}=0.1$ and $L=80$. Dashed grey lines show the fixed $\tilde{K}_{-}$ cut analyzed in Fig. 33. Obtained from MC simulations.
Fig. 33: Phase transition from disorder to swap in 3D in the general asymmetric case. (a) $R$ and (b) $\mathcal{L}$ as a function $\tilde{J}$ with $\tilde{K}_{+}=0.1$ and $\tilde{K}_{-}=0.3$, for different system sizes. The inset in panel a shows a snapshot of $\theta$ in a 2D slice of the 3D system (taken after steady state is reached), with $\tilde{J}=2$, $L=80$, and the same $\tilde{K}_{+}$ and $\tilde{K}_{-}$ values as in the main panel. Obtained from MC simulations.